A mass is oscillating with amplitude 𝐴 at the end of the spring. How far is this mass from the equilibrium if the elastic potential energy equals 𝛾 times the kinetic energy?



1
Expert's answer
2022-01-04T10:23:13-0500

total mechanical energy:

"\frac{kA^2}{2}=\frac{kx^2}{2}+\frac{mv^2}{2}"


we have:

"\frac{kx^2}{2}=\gamma\frac{mv^2}{2}"

then:

"\frac{kA^2}{2}=\frac{kx^2}{2}+\gamma\frac{kx^2}{2}"


"A^2=(1+\gamma)x^2"


displacement from the equilibrium:


"x=\frac{A}{\sqrt{1+\gamma}}"